OpenStudy (anonymous):

Determine whether the following polynomial is a perfect square trinomial: x^3 + 8x + 16 Yes or No

OpenStudy (anonymous):

lol but I'm 50% sure the answer is yes

OpenStudy (anonymous):

Lol what? xD

OpenStudy (anonymous):

@phi

OpenStudy (kirbykirby):

No, not if you have a degree-3 polynomial. Perfect square trinomial has the form: \((x+a)^2 = x^2+2ax+a^2\) or \((x-a)^2=x^2-2ax+a^2\)

OpenStudy (anonymous):

i think its no but how to find the sqrt of \(x^3\)

OpenStudy (phi):

if you have x^3 then it won't be

OpenStudy (anonymous):

but how do you know

OpenStudy (phi):

the square root of x^3 is \( x^{\frac{3}{2}} \) which is not "perfect" i.e. x

OpenStudy (anonymous):

yes keep going on

OpenStudy (kirbykirby):

You need to somehow get a form (smtg + smtg)^2 when you factor it

OpenStudy (phi):

they are assuming the square root will be a binomial of the form (x+a) where a is a number. or, if we relax the requirement and allow \[ (x^n + a) \] where n is an integer, we would get \[ (x^n + a)^2 = x^{2n}+ 2ax^n +a^2 \] and the highest exponent is even (2n will be an even number) the only way to get an exponent of 3 is to start with 3/2 and 3/2 is not integer.... and we insist on "simple" (x+a) or maybe (x^n +a) as the starting binomial... otherwise it's not perfect.

OpenStudy (anonymous):

ohhhhhhhhhhh, of course @phi \(x^3≠~perfect\)

OpenStudy (anonymous):

thx @phi so the answer is a no

OpenStudy (phi):

definitely NO